OpenStudy (dschneider2016):

Will give medal!! Please help!! Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (dschneider2016):

@mathstudent55 please help! Its super fast

OpenStudy (dschneider2016):

@dumbcow @LoveIt

OpenStudy (dumbcow):

treat it like a quadratic equation \[7u^2-14u +2 = -5\] solve for "u" then \[\sin x = u\] solve for x using inverse sine \[x = \sin^{-1} u\]

OpenStudy (anonymous):

I'm not too good at math, still in pre-algebra :(

OpenStudy (dschneider2016):

okay so would the next step be 7u^2−14u+7 u=1

OpenStudy (dschneider2016):

@dumbcow

OpenStudy (dschneider2016):

therefore sin x= 1 and to solve for x sin-1 (1) ?

OpenStudy (dumbcow):

good now plug u = 1 into inverse sine function or think what angles make sine = 1

OpenStudy (dschneider2016):

90°

OpenStudy (dumbcow):

correct

OpenStudy (dschneider2016):

okay so is that it?

OpenStudy (dumbcow):

depends on the domain given, for 0<x<360 thats it, one solution, x = 90 If they want all (infinite) solutions, then just add multiples of 360 to answer --> 90 + 360n

OpenStudy (dschneider2016):

So my answer is just 90 + 360n (on the assumption they want infinite solutions)?

OpenStudy (dschneider2016):

@dumbcow

OpenStudy (dumbcow):

yes

OpenStudy (dschneider2016):

Thanks so much! You're a life saver! @dumbcow

OpenStudy (dumbcow):

your welcome :)